OpenStudy (anonymous):

I don't believe all C's are created equal \[\int x^{2}cosmxdx\] u= x^2 dv=cosmx dx du=2x dx v=(1/m)sinmx \[\frac{x^{2}}{m}sin(mx)-\frac{2}{m}\int xsin(mx) dx+C\] u=x dv=sinmxdx du=dx v=-(1/m)cosmx \[-\frac{x}{m}cos(mx) + \frac{1}{m}\int cos(mx) dx \] \[(-\frac{x}{m}cos(mx) +\frac{1}{m^{2}}sin(mx) +C)\] \[\frac{x^{2}}{m}sin(mx)-\frac{2}{m}(-\frac{x}{m}cos(mx) +\frac{1}{m^{2}}sin(mx) +C)+C\] \[\frac{x^{2}}{m}sin(mx)+\frac{2x}{m^{2}}cos(mx) -\frac{2}{m^{3}}sin(mx) -\frac{2}{m}C+C\]

OpenStudy (anonymous):

how are my C's looking?

OpenStudy (anonymous):

why do you not use tabulars method

OpenStudy (anonymous):

x^2 goes to 0

OpenStudy (anonymous):

i was just thinking about that ...:)

OpenStudy (anonymous):

I have never used Tabulars method before, but I'm willing to learn

OpenStudy (anonymous):

i do believe there is a mistake as you shouldn't get that large of an integration

OpenStudy (anonymous):

it's a type of by parts method where your u goes to 0

OpenStudy (anonymous):

after so many derivatives

OpenStudy (anonymous):

so you make u = x^2 and dv = cos(mx)

OpenStudy (anonymous):

then you take as many derivatives as it takes to make ti got o0

OpenStudy (anonymous):